Two standard number cubes are rolled. What is the probability that the sum is 7 or an 11? The sum is a 2,3 or a 12? The sum is a 6 or 8? The sum is a &?
Answer by richard1234(7193)

Use this chart:
P(2) = P(12) = 1/36
P(3) = P(11) = 2/36
P(4) = P(10) = 3/36
P(5) = P(9) = 4/36
P(6) = P(8) = 5/36
P(7) = 6/36
For example, the first problem is just P(7) + P(11) = 8/36 = 2/9. Do the same for the other problems.